OpenStudy (anonymous):

4 − (x + 2) = 6 − 5(x + 4)

OpenStudy (anonymous):

Can you elaborate more

OpenStudy (mimi_x3):

I made a mistake lol; i read the question wrongly

OpenStudy (anonymous):

9x=22 would be the answer?

OpenStudy (anonymous):

oh... :( lol

OpenStudy (jiteshmeghwal9):

\[4-(x+2)=6-5x-20\]\[4-x-2=-5x-14\]\[2-x=-5x-14\]now take variables at one side & constants at one side\[-x+5x=-14+2\]\[4x=-12\]\[x=\frac{-12}{4}=-3\]

OpenStudy (mimi_x3):

\[ 4-(x+2) = 6- 5(x+4)\] 4 - x - 2 = 6 - 5x - 20 2-x = -5x -14 4x = -16 x = -4

OpenStudy (jiteshmeghwal9):

@Mimi_x3 u don't do the question like given

OpenStudy (mimi_x3):

@jiteshmeghwal9: you did not do it correctly :)

OpenStudy (jiteshmeghwal9):

ya! i didn't reverse the sign

OpenStudy (anonymous):

sooo its -4?

OpenStudy (mimi_x3):

Yes! Do you understand how to do it though? :)

OpenStudy (anonymous):

yes